OpenStudy (anonymous):

What is the y-intercept of the graph of the following linear equation? 5x-2y=20 A) -20 B) -10 C) 10 D) 20

OpenStudy (anonymous):

do you want the answer?

OpenStudy (anonymous):

No, i want an explanation. Then, after i get the explanation i want to try and see if i get the answer or not(:

OpenStudy (anonymous):

ok @freckles can you help him? @Monkey4224 can you help him?

OpenStudy (anonymous):

I am a her..

OpenStudy (anonymous):

opps sorry i mean can you help her?

OpenStudy (anonymous):

Okay..well it seems no one wants to help me xD

OpenStudy (anonymous):

im sorry but i cant help u either but ill give it a try

OpenStudy (anonymous):

Wait, is the answer 10?

OpenStudy (anonymous):

5x + 2y = 20 Subtract 5x from both sides 5x - 5x + 2y = 20 - 5x Simplify 2y = -5x + 20 Divide both sides by 2 2y / 2 = (-5x + 20) / 2 Simplify y = -2.5x + 10 ^ is that why it is -10?

OpenStudy (anonymous):

The correct answer is 10... which is what i said. https://answers.yahoo.com/question/index?qid=20090120124149AAOHmTS

OpenStudy (anonymous):

ok let me show you the work y-intercept = 20/-2=10/-1= -10.00000

OpenStudy (anonymous):

http://www.tiger-algebra.com/drill/5x-2y=20/ check this out